A generic reaction of aA———> products, gives a straight line plot of ln[A] versus time with a slope of -7.71x10^-3 Ms^-1. Assuming that the initial concentration of A is 0.135M, calculate the time in second,t, required for 25.6% of the original sample to react.


1
Expert's answer
2021-03-08T05:55:27-0500

The given generic reaction is a First Order reaction.


Slope = -k = "-7.71\times 10^{-3}"


Hence, k = "7.71\times 10^{-3}"


Initial concentration = a = 0.135M


We know in the first order reaction,


"t = \dfrac{2.303}{k}log\dfrac{a}{a-x}"


"t= \dfrac{2.303}{7.71\times 10^{-3}}log\dfrac{0.135}{0.135 - \dfrac{25.6}{100}\times 0.135}"


Hence, "t = 38.35" s